OpenStudy (anonymous):

3(x − 1) − (x + 5) = 2

mathslover (mathslover):

3(x-1)=? 3(x)-3(1)

mathslover (mathslover):

always remember that : a(b-c)=ab-ac hence 3(x-1)=? @maktaopve ?